Question 13 Using a computerized Inventory Management System, a Paint Supply Store franchise continuously monitors the invent can you show how to do this work
Answer #1

Answer 13:

d = Average Demand rate per time period

74

gallons per week

σ = Standard Deviation

8

gallons

L = Performance cycle, another term for total lead time

2.5

week

z for 99% target service level

2.326

T = time increment used for calculating standard deviation of demand

1

week

Safety stock =

z*(√L/T)*σ

Safety stock = z*(√L/T)*σ = (2.329)*(√2.5/1)*8 = 29.4

29

gallons